What is must be the velocity of the soccer ball to obtain a range of 30 m which kick is at an angle of 40 degrees?


1
Expert's answer
2021-09-20T16:32:12-0400

The range is

"R=\frac{v^2\sin(2\theta)}{g},\\\space\\
v=\sqrt{\frac{gR}{\sin(2 \theta)}}=17.3\text{ m/s}."